# Algebraic non-Hermitian skin effect and unified non-Bloch band theory in arbitrary dimensions

### v1

@@ IV Algebraic non-Hermitian skin effect in two dimensions
Through the established two-dimensional GBZ formula above, we can obtain the analytic OBC eigen wavefunctions constructed from the GBZ basis. Based on this, we demonstrate in this section that for the reciprocal non-Hermitian systems in higher dimensions, the skin modes generally follow the power-law localization. This universal algebraic skin effect is in sharp contrast to the exponential-law localization observed in conventional non-Hermitian skin effects.
Here, we prove that the reciprocal non-Hermitian systems lead to the algebraic skin effect in higher dimensions. A crucial point is that for a given energy $E$, when its 1D GFS curves in the complex momentum space extend through the Brillouin Zone, the resulting OBC wavefunction from the GFS basis exhibits power-law decay. In one dimension, the GFS for a given energy is 0D ($\dim{\text{GFS}}=0$), hence, this algebraic localization of skin effect is unique to dimensions greater than one. For demonstration, we use the model given by Eq. (21) and set the system parameters to $t_{x}=0,t_{y}=i,t_{xy}=1,u=-2i$. The non-Bloch Hamiltonian becomes
@@
which respects the reciprocity, $\mathcal{U}^{\dagger}H^{T}(\beta_{x},\beta_{y})\mathcal{U}=H(1/\beta_{x},1/\beta_{y})$, with the unitary matrix satisfying $\mathcal{U}\mathcal{U}^{\ast}=1$.
Figure 4:  The power-law localization of non-Hermitian skin effect. (a) The comparison between the PBC spectrum (the blue region) and OBC eigenvalues (the red dots), with the system size of $L_{x}=L_{y}=81$. (b) For the selected OBC eigenvalue $E_{0}=0.61-0.58i$ (the blue dot in (a)), there are four corresponding Fermi points $\bm{k}_{i}$ distributed in Brillouin zone, where $\det{[h(\bm{k}_{i})-E_{0}]}{=}0$. The projections of the GFS of $E_{0}$ on the $\beta_{x}$ and $\rho_{y}$ planes are shown in (c) and (d), respectively. The intersections between the GFS and the Brillouin zone (the black unit circle) are labeled by the purple and green dots, which correspond to the four Fermi points in (b). (e) The partial wavefunction that is constructed from the GFS inside the unit circle. (f) The layer density of the wavefunction exhibits power-law localization when it is away from the edge.
@@
When the Hamiltonian respects reciprocity, it exhibits the spectral property that OBC bulk eigenvalues belong to the Bloch spectrum, denoted as $E_{\text{OBC}}\in\sigma_{\text{PBC}}$. See a detailed proof in Appendix A. As a result, for a given OBC eigenvalue $E_{0}$ in reciprocal systems, there are always Bloch wave solutions $\bm{k}_{i}$ obtained by solving $\det[h(k_{x},k_{y})-E_{0}]=0$ where $h(k_{x},k_{y})$ denotes the Bloch Hamiltonian. These Bloch wave solutions, termed Fermi points, are related by reciprocity. In the above example with Hamiltonian given by Eq. (39), the Bloch spectrum is represented by the blue region in Fig. 4(a), and open-boundary eigenvalues for a system size of $L_{x}=L_{y}=81$ are indicated by red dots in Fig. 4(a). We select a generic OBC eigenvalue $E_{0}=0.61-0.58i$, denoted by the blue dot in Fig. 4(a). The distribution of the corresponding Fermi points in the Brillouin zone is illustrated in Fig. 4(b).
@@
As shown in Fig. 4(b), a pair of Fermi points $\bm{k}_{1}$ and $\bm{k}^{\prime}_{1}$ (or $\bm{k}_{2}$ and $\bm{k}^{\prime}_{2}$) have the same projection $k_{y,1}$ ($k_{y,2}$) on the $k_{y}$ axes. We can always let Bloch waves $k_{x,1}$ and $k_{x,3}$ form a standing wave, which is extended along the $x$ direction, and then transfer this standing wave along the $y$ direction, characterized by $k_{y,1}$. According to our standing-wave construction in Eq. (31), the OBC wavefunction of $E_{0}$ encompasses this standing-wave component: $\phi_{k_{y,1}}(x,y)=e^{ik_{y,1}y}(e^{ik_{x,1}x}-e^{ik_{x,3}x})$. The GFS for $E_{0}$ is obtained using the transfer matrix approach under the layer-$y$ basis, and its projections onto the $\beta_{x}$ and $\rho_{y}$ planes are shown in Figs. 4(c) and (d), respectively. The presence of Fermi points for $E_{0}$ means that its GFS must cross the Brillouin zone, as represented by the black unit circle in Figs. 4(c) and (d). These intersections, labeled by purple and green dots, correspond to the Fermi points in Fig. 4(b), specifically $\rho_{y,i}{=}e^{ik_{y,i}}$ and $\beta_{x,i}{=}e^{ik_{x,i}}$. In the continuum limit, the projection of the GFS on the $\rho_{y}$ plane forms 1D curves that cross the unit circle. Consequently, in reciprocal non-Hermitian systems, a generic OBC eigen wavefunction invariably encompasses Bloch standing wave components, characterized by $|\rho_{y}|=|\beta_{x}|=1$. The intersections between GFS and the Brillouin zone are guaranteed by the reciprocity of the Hamiltonian. These intersection points give rise to Bloch standing wave components of the OBC wavefunction. Next, we demonstrate that the presence of these Bloch wave components in the OBC wavefunction leads to the power-law localization.
@@
We focus on the intersection point $\rho_{y,1}$, which is relabeled by $\rho_{0}$ for notational clarity as shown in Fig. 4(d). In the continuum limit, the summation over $\rho_{y}$ points in the wavefunction of Eq. (31) becomes the integral over the $\rho_{y}$ curve. The contribution of wave components near the intersection $\rho_{0}$ is described by
@@
where $\delta\rho$ represents a piece of $\rho_{y}$ that is a small deviation from the $\rho_{0}$. Here, we study the localization of wavefunction at the bottom edge of the system. Therefore, we only select the $\rho_{y}$ components inside the unit circle and neglect the components outside. Hence, we choose $|\rho_{y}|\leq 1$ as the integral region. According to 2D GBZ condition in Eq. (30), we have $|\beta_{1}(\rho)|=|\beta_{2}(\rho)|=e^{\mu_{x}(\rho)}$. Consequently, the standing wave component in Eq. (40) becomes $\beta_{1}^{x}(\rho)-\beta_{2}^{x}(\rho)=e^{\mu_{x}(\rho)x}\sin{(\theta_{\rho}x)}$, where $\theta_{\rho}$ is the real-valued phase difference that depends on $\rho$. The intersection $\rho_{0}$ corresponds to $\beta_{x,1}$ and $\beta_{x,2}$ in the $\beta_{x}$ plane, where $|\beta_{x,1}|=|\beta_{x,2}|=1$. Near $\rho_{0}$ and within the unit circle, $|\beta_{x}(\rho)|\lesssim 1$ and $\mu_{x}(\rho)\lesssim 0$. The function $\mu_{x}(\rho)$ can be approximated as $\mu_{x}(\rho)=\mu_{x}(\rho_{0})+\mu^{\prime}_{x}(\rho_{0})\,\delta\rho+o(\delta\rho)\approx\mu_{x}(\rho_{0})=0$, where only the leading term is retained. Adopting this approximation, the Eq. (40) can be represented as
@@
We now define the layer density of the wavefunction to capture the decay law along the $y$ direction, expressed as
@@
where $|A_{0}|^{2}\neq 0$ is ensured when the open-boundary wavefunction includes the Bloch standing wave components, which holds true when the Hamiltonian respects reciprocity. So far, we have demonstrated that the presence of Bloch standing wave components leads to the universal power-law decay of ${r^{{-}1}}$ for the open-boundary wavefunctions, where $r$ is the distance from the open boundaries.
@@
The verification of power-law localization of wavefunction is shown in Figs. 4(e) and (f). To present the contribution of Bloch standing wave components more clear, we consider the partial OBC wavefunction that is constructed by the non-Bloch bases of GFS inside and on the unit circle, as denoted by the gray regions in Figs. 4(c) and (d). As shown in Fig. 4(e), this partial wavefunction is localized at the bottom edge $y$ due to localization length $\mu_{y}\leq 0$. Due to the reciprocity, the other part of the wavefunction constructed by the non-Bloch bases outside the unit circle will be localized at the top edge $y$ and is center-symmetric about the origin of the lattice. We calculate the layer density for this partial wavefunction, as shown by the black dots in Fig. 4(f), where the blue and red curves represent the exponential- and power-law fitting, respectively. It shows that the OBC wavefunction follows a power law tail of $y^{-1}$ when it moves away from the edge.
Remarkably, the power-law decay behavior is robust even with the finite-size system. For a system size of $L_{x}\times L_{y}$, there are $2L_{x}$ solutions for $\rho_{y}$. As shown in the inset of Fig. 4(d), it exhibits a minor gap between $\rho_{y}$ solutions and the unit circle. The gap scales with the length of $L_{x}$ and can be quantified as $\log{\Delta}\sim\alpha/L_{x}$, where $\alpha>0$ is a constant factor. This gap, determined solely by the length $L_{x}$ and independent of $L_{y}$, modifies the upper limit of the integral in Eq. (43) to $-\alpha/L_{x}$. Consequently, the average density with length of $L_{x}$ becomes:
$\mathcal{P}_{E_{0},\rho_{0},L_{x}}(y)\sim y^{-1}\,e^{-\alpha\,y/L_{x}},$ (45)
@@
which exponentially converges to a perfect power-law decay in Eq. (44) as $y$ increases. It concludes that the finite-size effects on decay behavior along the $y$ direction are entirely controlled by the system length $L_{x}$ in the $x$ direction. This is derived from the fact that a finite-size wavefunction is composed of $2L_{x}$ non-Bloch wave components and is independent of the length in the $y$ direction, hence establishing that the $L_{x}$-dependence of the power-decay behavior in the $y$ direction. This inter-dependency between the system length in the $x$ direction and the convergence rate to power-law decay in the $y$ direction is a characteristic feature of the algebraic non-Hermitian skin effect in a finite-size lattice system.
V The GBZ formula in generic higher dimensional systems
So far, we have established the GBZ formula in two dimensions using the layer transfer matrix method and derived the algebraic non-Hermitian skin effect. We now extend these formulas to generic higher dimensions.
@@ V The GBZ formula in generic higher dimensional systems
$({d=1})$ ${d-1}$ ${\dim{\text{GFS}}}$
area spectrum ${2d-1}$ ${2d-n}$ $d$ ${\dim{\text{GBZ}}}$
$({d\geq 2})$ ${2d-3}$ ${2d-n-2}$ ${d-2}$ ${\dim{\text{GFS}}}$
In two dimensions, we show that, unlike in one dimension, the number of standing wave conditions can be either one or two, producing a 3D or 2D GBZ manifold, respectively. In $d$ dimensions, the complex momenta span a $2d$-dimensional real space, and the number of standing wave conditions can be within the range of ${1\leq n\leq d}$. Correspondingly, the GBZ dimension is determined to be ${2d-n}$. Therefore, for a $d$-dimensional lattice system, the dimension of the GBZ manifold is in the range:
${d\leq\dim{\text{GBZ}}\leq 2d-1}.$ (46)
The GFS is related to the GBZ through the open-boundary spectrum, and they have the dimensionality relation: $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}$. Notably, when $d=1$, this inequality in Eq. (46) constrains the GBZ manifold to one dimension, and the open-boundary spectrum is restricted to forming some arcs in the complex energy plane, namely $\dim{E}=1$. Consequently, for a given OBC eigenvalue, the GFS consists of some points and has the dimensionality of $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}=0$. When $d\geq 2$, the case of GFS becomes more complicated due to the fact that the OBC eigenvalues can cover areas in the complex energy plane. In conclusion, the dimension of the GFS in $d\geq 2$ dimensions falls within the range:
For example, in $d=2$, the open-boundary spectrum covers a finite area in the complex energy plane, and the GBZ dimension could be either 2D or 3D. Correspondingly, the GFS is composed of 0D points or 1D curves. All these scenarios in two dimensions have been included in our previous discussions. The dimensions of GBZ and GFS in $d$-dimensional systems are concluded in Table 1.
According to the dimension counting in Eq. (46) and Eq. (47), in three dimensions (${d=3}$), the GBZ dimension could be $3$, $4$, or $5$, and the GFS dimension ranges from $1$ to $3$. In general cases, $\dim{\text{GBZ}}=5$ and $\dim{\text{GFS}}=3$. We now provide an example for this generic case.
For a three-dimensional non-Hermitian Hamiltonian, its open-boundary eigenvalues are typically complex-valued, and their number scales with the system size. In the most generic cases, these eigenvalues cover finite areas in the complex energy plane, that is, $\dim{E}=2$. For this case, we can extend the transfer matrix approach to three dimensions to obtain the GBZ formula and construct the analytic open-boundary wavefunctions. To demonstrate this generalization, we consider an example of a 3D non-Bloch Hamiltonian:
@@
where $t_{i=x,y,z}$ are the nearest neighbor hopping strengths along $x,y$, and $z$ directions, respectively. The term $t_{xyz}\neq 0$ represents hopping along the cube diagonal direction, effectively coupling the lower-dimensional systems in the $x,y$, and $z$ directions. One can first choose the transfer matrix basis. Without loss of generality, the Hamiltonian can be written in the planar-$xy$ basis and transferred along the $z$ direction. We assume the system size to be $L^{3}$, with $L$ being the system length. Similar to Eq. (24), a planar transfer matrix of dimension $2L^{2}\times 2L^{2}$ can be obtained. The eigenvalues of this transfer matrix, denoted as $\rho_{z}$, are allowed by the bulk Hamiltonian. In general, these $2L^{2}$ non-Bloch solutions of $\rho_{z}$ cover a finite region in the complex $\rho_{z}$ plane. For each fixed $\rho_{z_{0}}$, the 3D non-Bloch Hamiltonian Eq. (48) reduces to a 2D subsystem, $\mathcal{H}(\beta_{x},\beta_{y},\rho_{z_{0}})$. One can treat this 2D subsystem as a 2D system, and solve it using the layer transfer matrix approach once again in two dimensions. According to the established transfer matrix method for a 2D Hamiltonian, a 1D GFS curve in the $(\beta_{x},\rho_{y})$ space can be obtained, which varies with the choice of $\rho_{z_{0}}$. For each fixed $\rho_{z_{0}}$, 1D GFS curves in the $(\beta_{x},\rho_{y})$ for the 2D subsystem can be solved. As $\rho_{z}$ scans through the $2D$ region, a 3D GFS associated with a specific energy for this 3D Hamiltonian is ultimately obtained, indicating $\dim{\text{GFS}}=3$. Consequently, the GBZ for this Hamiltonian has the dimension $\dim{\text{GBZ}}=\dim{\text{GFS}}+\dim{E}=5$. For each fixed $\rho_{z_{0}}$ and $\rho_{y_{0}}$, the Hamiltonian reduces to 1D subsystem. Consequently, the standing wave condition along the $x$ direction can be enforced, which gives rise to the GBZ condition for this 3D system:
@@
Depending on the transfer matrix basis, the GBZ condition will change and is not unique, similar to the 2D cases discussed in subsection III.4. Similarly, the 3D OBC wavefunction can be constructed as follows: Let the non-Bloch waves first form standing waves along the $x$ direction; then transfer these standing waves along the $y$ direction to form planar components; each planar component is further transferred along the $z$ direction to generate a standing wave component. Therefore, the OBC wavefunctions for this model can be constructed with these standing wave components:

### v2

@@ IV Algebraic non-Hermitian skin effect in two dimensions
Using the above GFS framework, we can exactly construct open-boundary wavefunctions from the GFS basis. Based on this theoretical framework, this section presents a central result of this work: in dimensions higher than one, skin modes can exhibit quasi-long-range algebraic localization, with wavefunction amplitudes decaying as a power law away from the boundary. This behavior contrasts sharply with the conventional expectation of exponential localization in the non-Hermitian skin effect.
The emergence of the algebraic skin effect hinges on the following key mechanism. In two dimensions, an OBC eigenstate consists of a continuum of non-Bloch wave components determined by the GFS curves. The wavefunction exhibits algebraic localization if and only if these components include an extended Bloch wave with a diverging localization length. The necessary and sufficient conditions for the appearance of the algebraic skin effect can be formally stated as follows: (i) $\dim(\text{GFS})\geq 1$; (ii) the GFS curve contains real momentum components, i.e., it intersects the Brillouin zone with nonzero superposition weights. These conditions restrict the algebraic skin effect to systems with spatial dimension higher than one. In principle, they can be satisfied in both reciprocal and non-reciprocal systems, implying that the algebraic skin effect may arise in any non-Hermitian system beyond one dimension.
Under these two conditions, we further deduce that in two or higher dimensions, reciprocal non-Hermitian skin effects, such as the geometry-dependent skin effect [81], generally manifest as algebraic localization under generic boundary geometries. In the following, we illustrate this sufficient condition using a reciprocal skin effect model and present a non-reciprocal example that exhibits algebraic skin modes in Appendix C.
[…]
For demonstration, we consider the reciprocal model given by Eq. (21) with system parameters set to $t_{x}=0$, $t_{y}=i$, $t_{xy}=1$, and u=2iu=-2i. The non-Bloch Hamiltonian becomes
@@
which respects the reciprocity $H^{T}(\beta_{x},\beta_{y})=H(1/\beta_{x},1/\beta_{y})$. When the Hamiltonian is reciprocal, it exhibits the spectral property that all bulk eigenvalues under OBCs lie within the Bloch spectrum, i.e., $E_{\text{OBC}}\in\sigma_{\text{PBC}}$. A direct proof is provided in Appendix D. As a result, for a given OBC eigenvalue $E_{0}$ in reciprocal systems, there are always Bloch-wave solutions $\bm{k}_{i}$ obtained by solving $\det[h(k_{x},k_{y})-E_{0}]=0$ where $h(k_{x},k_{y})$ denotes the Bloch Hamiltonian. These Bloch wave solutions, termed Fermi points, necessarily lie on the GFS curves in reciprocal systems.
@@
For the Hamiltonian given by Eq. (39), the Bloch spectrum is represented by the blue region in Fig. 5(a), and open-boundary eigenvalues for a system size of $L_{x}=L_{y}=81$ are indicated by red dots in Fig. 5(a). We select a generic OBC eigenvalue $E_{0}=0.61-0.58i$, denoted by the blue dot in Fig. 5(a). The distribution of the corresponding Fermi points in the Brillouin zone is illustrated in Fig. 5(b). As shown in Fig. 5(b), a pair of Fermi points, $\bm{k}_{1}$ and $\bm{k}^{\prime}_{1}$ (or $\bm{k}_{2}$ and $\bm{k}^{\prime}_{2}$), project to the same point $k_{y,1}$ ($k_{y,2}$) on the $k_{y}$ axes. Therefore, we can let Bloch waves $k_{x,1}$ and $k_{x,3}$ form a standing wave in the $x$ direction, which then propagates along the $y$ direction characterized by $k_{y,1}$. This forms one Bloch standing-wave component of the OBC wavefunction. Therefore, these Fermi points are components on the GFS. The discretized GFS for $E_{0}$ is obtained using the transfer matrix approach within the $x$-standing wave condition. The projections of GFS onto the complex $\beta_{x}$ and $\rho_{y}$ planes are shown in Figs. 5(c) and (d), respectively. The presence of Fermi points for $E_{0}$ means that its GFS must intersect the Brillouin zone, represented by the black unit circle in Figs. 5(c) and (d). These intersections, marked by purple and green dots, correspond to the Fermi points in Fig. 5(b), specifically $\rho_{y,i}{=}e^{ik_{y,i}}$ and $\beta_{x,i}{=}e^{ik_{x,i}}$. In the thermodynamic limit, the GFS on the $\rho_{y}$ plane forms 1D continuous curves that crosses the unit circle. Consequently, in reciprocal non-Hermitian systems, a generic OBC eigenstate always includes Bloch standing-wave components characterized by $|\rho_{y}|=|\beta_{x}|=1$. The intersections between GFS and the Brillouin zone are guaranteed by the reciprocity of the Hamiltonian. Next, we demonstrate that the presence of these Bloch wave components in the OBC wavefunction leads to a power-law localization.
@@
We focus on the intersection point $\rho_{y,1}$, which is relabeled by $\rho_{0}$ for notational clarity as shown in Fig. 5(d). In the continuum limit, the summation over $\rho_{y}$ points in the wavefunction of Eq. (31) becomes the integral over the $\rho_{y}$ curve. The contribution of wave components near the intersection $\rho_{0}$ is given by
@@
[…]
We now define the layer density of the wavefunction to capture the decay behavior along the $y$ direction, expressed as
@@
where $|A_{0}|^{2}\neq 0$ means that the open-boundary wavefunction includes the Bloch standing-wave components, which holds true when the Hamiltonian respects reciprocity. See more detailed discussion in Appendix E. So far, we have demonstrated that the presence of Bloch standing-wave components leads to the power-law decay $r^{{-}\alpha}$} ($\alpha\geq 1$) for the open-boundary wavefunctions, where $r$ is the distance from the open boundaries and the exponent $\alpha$ is determined by the structure of the GFS near the unit circle.
@@
The power-law localization of the wavefunction is verified in Figs. 5(e) and (f). To clearly demonstrate the algebraic decay, we examine the OBC wavefunction components near the bottom edge, which are constructed from the GFS bases inside and on the unit circle, as indicated by the gray regions in Figs. 5(c) and (d). As shown in Fig. 5(e), this partial wavefunction is localized at the bottom edge due to a localization exponent $\mu_{y}\leq 0$. By reciprocity, the remaining components, constructed from non-Bloch basis elements outside the unit circle, are localized at the top edge. We calculate the layer density for this partial wavefunction, shown as black dots in Fig. 5(f), with the blue and red curves representing exponential and power-law fittings, respectively. The results clearly reveal that the OBC wavefunction exhibits a power-law tail, decaying as $y^{-\alpha}$ away from the bottom edge.
[…]
$\mathcal{P}_{E_{0},\rho_{0},L_{x}}(y)\sim y^{-1}\,e^{-c\,y/L_{x}},$ (45)
@@
which exponentially converges to a perfect power-law decay in Eq. (44) as $y$ increases. The finite-size effects on decay behavior along the $y$ direction are determined by the system length $L_{x}$ in the $x$ direction. This follows from the fact that a finite-size wavefunction is composed of $2L_{x}$ non-Bloch wave components and is independent of the length in the $y$ direction.
V The robust algebraic localization with boundary disorders
Figure 6:  The robustness of algebraic skin effect with random boundary disorder. The spectrum and wavefunction corresponding to $E_{0}=0.61-0.52i$ without and with the presence of boundary disorder are shown in (a)(b) and (c)(d), respectively. To simulate the boundary disorder, we use onsite potentials with strengths randomly distributed in the range of $[0,W]$ at the layers $y=1$ and $y=L_{y}$. (c) and (d) are results for $W=1/100$. In the inset of (e), we plot the partial wavefunction constructed from GFS basis inside the unit circle. The black dots in (e) represent layer density decaying along the $y$ direction, with blue and red curves showing linear and power-law fittings ($y^{-\alpha}$ with $\alpha\geq 1$), respectively. (f) shows a log-log plot for different strengths of boundary disorder $W$. The relation between disorder strength $W$ and power-law decay index $\alpha$ is shown in the inset of (f). Disorder strengths are calculated for $W=0$ (blue color), $0.01$ (red), $0.1$ (pink), $0.2$ (green), $0.3$ (cyan), $0.5$ (purple) and $W=1$ (yellow).
The algebraic decay behavior in the higher-dimensional skin effect is robust, even in the presence of random open-boundary disorders. In this section, we numerically verify the robustness of the algebraic skin effect against boundary disorder. The main results are illustrated in Fig. 6. As shown in Fig.6(f), even weak boundary disorder, by breaking reciprocity, rapidly enhances the power-law decay. As the disorder strength increases from zero, the decay exponent quickly saturates to a steady value (Fig.6(f)), indicating that the algebraic decay is not only amplified but also stabilized by the disorder. These results confirm that the algebraic skin effect persists and remains robust even under strong boundary disorder.
We use the reciprocal example to demonstrate the robustness of the algebraic skin effect, whose Hamiltonian is given by Eq. (39). The Hamiltonian under square OBC geometry respects reciprocity. Once the random boundary disorder is introduced, the reciprocity is broken at the boundary. According to our GFS framework, the GFS curve is solely determined by the bulk Hamiltonian, and the specific boundary conditions determine the superposition coefficients of these GFS basis. Therefore, even with the random boundary disorder, the GFS curve is the same as that without boundary disorder. In Section IV, we have demonstrated that the GFS curve for this reciprocal Hamiltonian intersects the Brillouin zone. In addition, the random boundary disorder cannot diminish the Bloch standing-wave components. These two reasons guarantee the robustness of the existence of algebraic skin effect.
In the numerical simulation, we set the boundary disorder to be on-site potentials with strengths randomly distributed in the range of $[0,W]$ at the boundary layers $y=1$ and $y=L_{y}$. For the reciprocal Hamiltonian in Eq. (39), the OBC spectra without and with boundary disorder of strength $W=1/100$ are shown in Figs. 6(a) and (c), respectively, with the system size $L_{x}=L_{y}=61$. The spectrum density exhibits extreme sensitivity to the weak boundary disorder, which is due to the reciprocity breaking induced by the random boundary disorder. Correspondingly, the shapes of wavefunctions of OBC eigenvalue $E_{0}\approx 0.61-0.52i$ are largely modified by the boundary disorder, as shown in Figs. 6(b) and (d), respectively. Despite the large modification in the wavefunction, its power-law decay behavior remain robust even with the presence of boundary disorder. In Fig. 6(e), we plot the layer density (the black dots) for the partial wavefunction constructed by the GFS basis inside the unit circle, which follows a similar approach as Fig. 5(e) but has different superposition coefficients modified by the boundary disorder. The blue line and the red curve in Fig. 6(e) refers to the exponential and power-law fittings respectively, which clearly shows the power-law decay of the layer density along the $y$ direction, i.e., $\mathcal{P}_{E_{0}}(y)\propto y^{-\alpha}$ where the $\alpha$ is the power-law decay index. A smaller value of $\alpha$ indicates a slower power-law decay away from the boundary. In Fig. 6(f), we show the log-log plot of the layer density for different strengths $W$ of boundary disorder. Disorder strengths are calculated for $W=0,0.01,0.1,0.2,0.3,0.5$, and $1$, marked in different colors. The straight lines in the log-log plot are linear fittings indicating the power law decay behavior. Notably, as the boundary disorder strength increases from zero, which corresponds to the blue fitting line and blue dot in the inset of Fig. 6(f), the power-law exponent $\alpha$ saturates exponentially to a steady value, as shown in the inset of Fig. 6(f). The simulation results indicate the role of boundary disorder. The presence of boundary disorder breaks reciprocity, making the GFS components mix sufficiently. Consequently, the power-law decay behavior is stabilized by the boundary disorder, which demonstrates the robustness of the algebraic skin effect even with the presence of boundary disorder.
VI The GFS formula for generic higher dimensional lattice systems
Thus far, we have established the GFS formula in two dimensions, based on the standing-wave construction illustrated in Fig. 1, which allows us to construct the OBC wavefunctions. Building on this foundation, we derived the quasi-long-range algebraic decay characteristic of the 2D non-Hermitian skin effect. We now extend these conclusions to generic higher dimensions.
In two dimensions, we show that, unlike in one dimension, the number of standing wave conditions can be either one or two, leading to a 3D or 2D GBZ manifold, respectively. In $d$ dimensions, the complex momenta span a $2d$-dimensional real space, and the number of standing wave conditions can be within the range of $1\leq n\leq d$}. Correspondingly, the GBZ dimension is determined to be 2dn{2d-n}. Therefore, for a $d$-dimensional lattice system, the dimension of the GBZ manifold is in the range:
$d\leq\dim{\text{GBZ}}\leq 2d-1$}. (46)
The GFS is submanifold embedded in the GBZ for a fixed open-boundary eigenenergy, and they follow the dimensionality relation: $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}$. Notably, when $d=1$, this inequality in Eq. (46) constrains the GBZ manifold to one dimension, and the open-boundary spectrum is restricted to forming some arcs in the complex energy plane, namely $\dim{E}=1$. Consequently, for a given OBC eigenvalue, the GFS consists of some points and has the dimensionality of $\dim{\text{GFS}}=\dim{\text{GBZ}}-\dim{E}=0$. When $d\geq 2$, the case of GFS becomes more complicated due to the fact that the OBC eigenvalues can cover areas in the complex energy plane. In conclusion, the dimension of the GFS in $d\geq 2$ dimensions falls within the range:
For example, in $d=2$, the open-boundary spectrum covers a finite area in the complex energy plane, and the GBZ dimension could be either 2D or 3D. Correspondingly, the GFS is composed of 0D points or 1D curves. All these scenarios in two dimensions have been discussed in our previous discussions. The dimensions of GBZ and GFS in $d$-dimensional systems are summarized in Table 1.
@@ VI The GFS formula for generic higher dimensional lattice systems
(d=1)({d=1}) d1{d-1} $\dim{\text{GFS}}$}
area spectrum 2d1{2d-1} 2dn{2d-n} $d$ $\dim{\text{GBZ}}$}
$({d\geq 2})$ 2d3{2d-3} 2dn2{2d-n-2} d2{d-2} $\dim{\text{GFS}}$}
According to the dimension counting in Eq. (46) and Eq. (47), in three dimensions (d=3{d=3}), the GBZ dimension could be $3$, 44, or $5$, accordingly, the GFS dimension ranges from $1$ to $3$. For most generic case, $\dim{\text{GBZ}}=5$ and $\dim{\text{GFS}}=3$. We now provide an example for this generic case.
For a three-dimensional non-Hermitian Hamiltonian, its open-boundary eigenvalues are typically complex-valued, and their number scales with the system size. In the most generic cases, these eigenvalues cover finite areas in the complex energy plane, that is, $\dim{E}=2$. For this case, we can extend the transfer matrix approach to three dimensions to obtain the GFS formula and construct the analytic open-boundary wavefunctions. To demonstrate this generalization, we consider an example of a 3D non-Bloch Hamiltonian:
@@
where $t_{i=x,y,z}$ are the nearest neighbor hopping strengths along $x,y$, and $z$ directions, respectively. The term $t_{xyz}\neq 0$ represents hopping along the cube diagonal direction, effectively coupling the lower-dimensional systems in the $x,y$, and $z$ directions. One can first choose the transfer matrix basis. Without loss of generality, the Hamiltonian can be written in the planar-$xy$ basis and transferred along the $z$ direction. We assume the system size to be $L^{3}$, with $L$ being the system length. Similar to Eq. (24), a planar transfer matrix of dimension $2L^{2}\times 2L^{2}$ can be obtained. The eigenvalues of this transfer matrix, denoted as $\rho_{z}$, are allowed by the bulk Hamiltonian. In general, these $2L^{2}$ non-Bloch solutions of $\rho_{z}$ cover a finite region in the complex $\rho_{z}$ plane. For each fixed $\rho_{z_{0}}$, the 3D non-Bloch Hamiltonian Eq. (48) reduces to a 2D subsystem, $\mathcal{H}(\beta_{x},\beta_{y},\rho_{z_{0}})$. One can treat this 2D subsystem as a 2D system, and solve it using the layer transfer matrix approach once again in two dimensions. According to the established transfer matrix method for a 2D Hamiltonian, a 1D GFS curve in the $(\beta_{x},\rho_{y})$ space can be obtained, which varies with the choice of $\rho_{z_{0}}$. For each fixed $\rho_{z_{0}}$, 1D GFS curves in the $(\beta_{x},\rho_{y})$ for the 2D subsystem can be solved. As $\rho_{z}$ scans through the 2D region of the transfer-matrix eigenvalues, a 3D GFS for the fixed energy is obtained in this 3D system, indicating $\dim{\text{GFS}}=3$. Consequently, the GBZ for this Hamiltonian has the dimension $\dim{\text{GBZ}}=\dim{\text{GFS}}+\dim{E}=5$.
For each fixed $\rho_{z_{0}}$ and $\rho_{y_{0}}$, the Hamiltonian reduces to 1D subsystem. Consequently, the standing wave condition along the $x$ direction can be enforced, which gives rise to the GFS condition for this 3D system:
@@
Depending on the transfer matrix basis, the GFS condition will change and is not unique, similar to the 2D cases discussed in subsection III.4. Similarly, the 3D OBC wavefunction can be constructed as follows: Let the non-Bloch waves first form standing waves along the $x$ direction; then transfer these standing waves along the $y$ direction to form planar components; each planar component is further transferred along the $z$ direction to generate a standing wave component. Therefore, the OBC wavefunctions for this model can be constructed with these standing wave components: